# A Convex-Inspired Neural Construction for Structured and Generalizable Nonlinear Model Reduction

### v1

@@ 4.1. Motivation
In nonlinear model reduction, the displacement field is expressed as a nonlinear mapping $\mathbf{u}=f(\mathbf{q})$ from reduced coordinates $\mathbf{q}\in\mathbb{R}^{k}$. Our goal is to design a nonlinear reduced-order structure that generalizes well while maintaining a compact representation of the reduced space.
A common approach is to introduce regularization terms that promote smoothness or other desirable properties in the latent space (Sharp et al., 2023; Lyu et al., 2024; Liu et al., 2022). While effective within the sampled training domain, these methods impose regularity only implicitly through the loss function and typically require extensive tuning to balance against task-specific objectives. Moreover, such regularizers act only on sampled regions of the latent space; because this space is generally unbounded, it is infeasible to enforce consistent behavior over the entire domain during training.
To address these limitations, we regularize the nonlinear latent space through the design of the mapping $f:\mathbb{R}^{k}\rightarrow\mathbb{R}^{N}$ itself, rather than through explicit loss terms. To avoid oscillatory behavior and improve stability under out-of-distribution inputs, we incorporate a convex-inspired structure into $f$. Convexity acts as an implicit regularizer, shaping the function over the entire latent domain. In addition, we introduce a symmetric (odd-function) architectural component to ensure consistent responses under loadings with inverted directions—cases often absent during training but common at test time.
We combine these ideas by embedding ICNN elements and symmetry constraints within the architecture of $f$, producing a nonlinear reduced-order representation with improved robustness and generalization. Figure 4 provides an overview of this formulation.
Figure 6. Effect of the odd-symmetric construction. Our odd-function formulation models the intrinsic symmetry of the system, enabling better generalization to deformations with opposite displacements. We train the model on a single deformation trajectory and visualize the resulting displacements from both the training latent codes and their inverses. Without the odd-function term (left), the model produces plausible deformations for training codes but exhibits higher variance and artifacts under latent inversion. With the odd-symmetric formulation (right), the deformations remain consistent across both original and inverted latent codes, demonstrating improved symmetry and generalization.
4.1.1. Convex-inspired Regularization in Reduced-space
Instead of learning a nonlinear mapping directly from the reduced coordinates $\mathbf{q}\in\mathbb{R}^{k}$ to the full displacement field $\mathbf{u}\in\mathbb{R}^{N}$, we introduce an intermediate latent space and apply convexity only to this stage. Specifically, we construct a mapping
$\mathbf{z}=f_{\text{convex}}(\mathbf{q}),\qquad f_{\text{convex}}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{r},\quad k<r\ll N,$
where $f_{\text{convex}}$ is implemented using ICNN-inspired components. The full displacement is then obtained through a linear basis:
$\mathbf{u}=\mathbf{W}\,\mathbf{z},\qquad\mathbf{W}\in\mathbb{R}^{N\times r}.$
Similar staged nonlinear–linear constructions have been explored in (Fulton et al., 2019).
First, convexity governs only the mapping $\mathbf{q}\mapsto\mathbf{z}$, fixing the sign of its second derivative and thereby discouraging oscillatory behavior throughout the entire latent space, including unsampled regions. At the same time, the unconstrained linear map $W$ provides flexibility, effectively relaxing the latent manifold while preserving the curvature monotonicity enforced by convexity.
Second, since $W$ is linear and sign-unconstrained, it can be initialized using PCA or other data-driven bases, providing a strong starting point that accelerates training and improves final reconstruction accuracy.
As seen in Figure 5, this formulation achieves faster convergence and a lower final loss compared to applying input convexity directly to the $\mathbf{q}\!\rightarrow\!\mathbf{u}$ mapping.
4.1.2. Odd Function Formulation
For reduced-space deformable simulation, especially in near-isometric settings, it is often desirable for the mapping to exhibit odd symmetry. This ensures that opposite reduced coordinates produce displacements of equal magnitude and opposite direction, reflecting the symmetric response of many physical systems. Odd symmetry also improves generalization: deformations corresponding to unseen loadings can be extrapolated through mirrored configurations of the training data. This structure is also consistent with classical linear model reduction, where linear bases naturally induce an odd mapping from reduced coordinates to full-space displacements. A practical benefit is that oddness guarantees zero displacement at $\mathbf{q}=0$, eliminating the need to explicitly encode the rest state.
To impose this property by construction, we evaluate the convex component $f_{\text{convex}}$ on both $\mathbf{q}$ and $\mathbf{q}$ and take their difference. The final mapping is therefore defined as
(6) $f(\mathbf{q})=\mathbf{W}\,f_{\text{convex}}(\mathbf{q})-\mathbf{W}\,f_{\text{convex}}(-\mathbf{q}),$
which immediately yields
$f(-\mathbf{q})=-\,f(\mathbf{q}),$
ensuring an odd mapping consistent with the physical symmetry of many loading conditions.
As illustrated in Figure  6, this odd formulation promotes generalization beyond the training range by symmetrically extending learned deformations to inverse directions in the reduced space.
4.2. Implementation
@@ 5. Results
We test our method across various settings. For all cases, the training data were generated using a full-space FEM solver with stable Neo-Hookean elastic energy (Smith et al., 2018). All neural networks were trained using the Adam optimizer (Kingma and Ba, 2015) for 50,000 epochs with a learning rate of $1\times 10^{-4}$. Unless otherwise noted, all neural networks employ PCA-based initialization for their first and last layers; the only exception is the camel-collapse experiment, which uses standard initialization.
@@ 5.1. Generalization on Unseen Loading
We evaluate our model’s generalization capability under two distinct out-of-distribution (OOD) loading scenarios: opposite-direction loading on the Frog (Fig. 7) and larger-magnitude loading on the Armadillo (Fig. 8). In both cases, the data are the full-space simulation data within a limited force range and tested beyond the training distribution to assess robustness.
@@
For the Frog (Fig. 7) , the training data consist solely of compressive deformations generated by downward loading, whereas the test scenario applies an equal-magnitude force in the opposite (tensile) direction. This setup explicitly probes the model’s ability to extrapolate beyond the training regime and recover physically consistent behavior under reversed loading conditions. During evaluation, the momentum term is removed and the system is relaxed toward a low-energy configuration under tension. The linear basis model reduction preserves a coherent, physically meaningful stretching pattern across iterations. The vanilla autoencoder (AE) remains numerically stable but fails to produce a plausible stretching deformation, often generating distorted shapes inconsistent with the expected tensile behavior. In contrast, our symmetric odd-function formulation achieves shape quality comparable to the linear basis, faithfully capturing the intended stretch. Moreover, it converges to a low-energy configuration faster, highlighting its improved generalization under unseen loading directions.
@@
For the Armadillo (Fig. 8), the training dataset covers moderate loading magnitudes, while the test scenario applies substantially stronger forces that exceed the training range. This experiment evaluates the framework’s ability to generalize to high-strain regimes where nonlinear effects dominate. At the beginning of the deformation evolution, both the linear basis and the vanilla autoencoder (AE) produce visibly distorted shapes under the strong load. As the relaxation proceeds, all methods gradually recover to a plausible configuration. However, the energy plot in Fig. 8 reveals distinct convergence behavior: our symmetric and convex formulation achieves markedly faster energy reduction than the linear baseline, while the vanilla AE plateaus at a higher energy level and fails to reach the minimal state. These results highlight our formulation’s efficiency and robustness when extrapolating to unseen, high-magnitude loading conditions.

### v2

@@ 4.1. Motivation
In nonlinear model reduction, the displacement field is represented as a nonlinear mapping $\mathbf{u}=f(\mathbf{q})$ from reduced coordinates $\mathbf{q}\in\mathbb{R}^{k}$. Our objective is to design a reduced-order structure that generalizes beyond the training distribution, particularly to larger load magnitudes and reversed force directions, while maintaining a compact latent representation.
Existing approaches typically rely on loss-based regularizers to encourage smoothness or stability in the latent space (Sharp et al., 2023; Lyu et al., 2024; Liu et al., 2022). Although effective within the sampled domain, such regularization acts only implicitly through the objective function and requires careful tuning against task-specific losses. Moreover, these penalties constrain behavior only at sampled points; since the latent space is unbounded, they provide no guarantee of consistent responses in unsampled regions, where extrapolation is most critical.
We instead embed the desired inductive bias directly into the mapping $f:\mathbb{R}^{k}\rightarrow\mathbb{R}^{N}$. Our formulation adopts an odd-symmetric Difference-of-Convex (odd-DC) structure, expressing $f$ as the difference of two convex functions. Convexity supplies a global regularity prior that discourages oscillations across the entire latent space, improving extrapolation to unseen magnitudes. Antisymmetrizing the convex primitive introduces an odd mapping, enabling consistent responses under inverse loading directions, a property not attainable with convexity alone.
This DC construction relaxes overly restrictive linear assumptions such as PCA while retaining structured smoothness, yielding a nonlinear decoder that responds more coherently on out-of-distribution reduced states. Figure 2 summarizes the proposed formulation.
4.2. Regularizing the Reduced Space
Rather than imposing structure directly on the full displacement mapping, we apply regularization in an intermediate reduced space. We define a latent representation
(5) $\mathbf{z}=f(\mathbf{q}),\qquad f:\mathbb{R}^{k}\rightarrow\mathbb{R}^{r},\quad k<r\ll N,$
and reconstruct the displacement through a linear trainable basis
(6) $\mathbf{u}=\mathbf{W}\,\mathbf{z},\qquad\mathbf{W}\in\mathbb{R}^{N\times r}.$
Similar nonlinear–linear staged constructions were explored in (Fulton et al., 2019) for their efficiency and stability benefits. In our formulation, this separation localizes structural constraints to the latent mapping while leaving the reconstruction layer $\mathbf{W}$ unaffected. As shown in Figure 4, this organization also leads to faster convergence in our setting. The decoupling allows the nonlinear stage to be regularized without over-restricting the final displacement representation and keeps $\mathbf{W}$ replaceable for later extensions such as neural-field bases.
4.3. Odd-Symmetric Difference-of-Convex Construction
To improve generalization to both larger load magnitudes and inverse force directions, we structure the latent mapping using an odd difference-of-convex composition,
(7) $f(\mathbf{q})=f_{\text{convex}}(\mathbf{q})-f_{\text{convex}}(-\mathbf{q}),$
which satisfies $f(-\mathbf{q})=-f(\mathbf{q})$ by design. The nonconvexity of the overall mapping arises solely from the difference of two convex evaluations, while the only trainable nonlinear component remains the convex primitive $f_{\text{convex}}$. This antisymmetrization is a fixed transformation applied at evaluation and therefore does not enlarge the optimization search space. As illustrated in Figure 3, the odd-DC formulation extends learned deformations coherently to inverse directions, promoting consistent responses outside the training range.
The above construction requires a parameterization of $f_{\text{convex}}$ that remains convex for all inputs; generic neural networks cannot guarantee this property. Consequently, a model class with built-in convexity is essential to realize the proposed structural prior.
4.4. Input Convex Network as the Convex Primitive
We construct the convex primitive function $f_{\text{convex}}$ using an Input Convex Neural Network (ICNN) (Amos et al., 2017), which guarantees convexity of the output with respect to the input by design.
As shown in Figure 5, an ICNN defines hidden activations $\mathbf{h}_{0},\dots,\mathbf{h}_{m}$ via
(8) $\mathbf{h}_{i+1}=g_{i}\!\left(\mathbf{W}_{i}^{(h)}\mathbf{h}_{i}+\mathbf{W}_{i}^{(q)}\mathbf{q}+\mathbf{b}_{i}\right),\quad i=0,\dots,m{-}1,$
with $f_{\text{convex}}(\mathbf{q})=\mathbf{h}_{m}$. Convexity is ensured by enforcing $\mathbf{W}_{i}^{(h)}\!\geq\!0$ for $i\!\geq\!1$ and using convex, non-decreasing activation functions (e.g., ReLU or softplus); see (Amos et al., 2017) for details.
Figure 6 illustrates this property on a didactic 2D example. Both an ICNN and a standard MLP are trained to fit the convex function $z=x^{2}+y^{2}$ using samples within a radius-2 disk. While the MLP exhibits oscillatory behavior outside the training region, the ICNN extrapolates smoothly due to its convex structure.
By imposing convexity only on the latent mapping $\mathbf{q}\mapsto\mathbf{z}$, the ICNN suppresses unwanted oscillations globally, including in regions without training data. The subsequent linear reconstruction via $\mathbf{W}$ (see Equation 5) retains full expressiveness, ensuring that convexity regularizes only the reduced space without limiting the final displacement field.
4.5. Training
We adopt a data-driven approach using a dataset of observed configurations $Q=\{\mathbf{u}_{i}\in\mathbb{R}^{N}\}$. At inference time, the reduced model consists solely of the decoder mapping $\mathbf{u}=\mathbf{W}f(\mathbf{q})$. During training, however, the latent coordinates $\mathbf{q}$ corresponding to each $\mathbf{u}_{i}$ are unknown, and an encoder $E:\mathbb{R}^{N}\!\rightarrow\!\mathbb{R}^{k}$ is introduced to infer them. We therefore train an autoencoder whose decoder is the proposed structured mapping $f$, while the encoder $E$ is used only to obtain training-time estimates $\mathbf{q}_{i}=E(\mathbf{u}_{i})$. The form of $E$ depends on the representation and is described separately for mesh-based and continuous reductions.
4.5.1. Mesh-based Model Reduction
For mesh-based reductions, the encoder $E$ is parameterized as an MLP that maps the $N$-dimensional displacement vector to a $k$-dimensional latent code. Following (Fulton et al., 2019), we initialize the first linear layer of $E$ and the basis $\mathbf{W}$ using PCA. The components $E$, $f$, and $\mathbf{W}$ are jointly optimized with the reconstruction loss
(9) $\mathcal{L}_{\text{mesh}}=\sum_{\mathbf{u}_{i}\in Q}\big\|\mathbf{W}f(E(\mathbf{u}_{i}))-\mathbf{u}_{i}\big\|^{2}.$
Figure 5. The convex component in odd-DC is built upon the input-convex neural network (ICNN), in which the output is, by construction, convex with respect to the input. This convexity is ensured through non-negative weights and non-decreasing activation functions.
Figure 6. Convexity-enforcing architectures enable smooth extrapolation. An input-convex network and a standard MLP are trained on samples of $z=x^{2}+y^{2}$ within a disk of radius 2. The convex model preserves the global convex structure of the function and generalizes smoothly outside the training region, whereas the unconstrained MLP exhibits non-smooth artifacts when extrapolating.
4.5.2. Continuous Model Reduction
For continuous model reduction, the mesh discretization may vary between training and inference: the number of vertices $N$ and their ordering are not fixed. This prevents parameterizing the encoder $E$ and basis $\mathbf{W}$ as fixed-size matrices. Importantly, the dimension of the reduced space remains constant, so the latent mapping $f(\mathbf{q}):\mathbb{R}^{k}\!\rightarrow\!\mathbb{R}^{r},\;k<r\ll N$ from Eq. 5 is still applicable.
To achieve discretization-agnostic reconstruction, we replace the trainable matrix $\mathbf{W}$ with a continuous basis function following (Chang et al., 2023). The basis is represented as a neural field $\mathbf{W}(\mathbf{x}_{0})$ that maps the rest position of any vertex $i$, $\mathbf{x}_{0}^{i}$, to a local basis $\mathbf{W}_{i}\in\mathbb{R}^{d\times r}$, where $d\!\in\!\{2,3\}$ is the spatial dimension. Because the latent dimension $r$ is fixed, the full-space displacement for an arbitrary set of vertices can be reconstructed as
(10) $\mathbf{u}(\mathbf{x}_{0})=\mathbf{W}(\mathbf{x}_{0})\,f(\mathbf{q}),$
independent of mesh resolution or ordering.
The remaining question is how to construct an encoder compatible with this discretization-agnostic representation. Because the basis dimension $r$ is fixed regardless of the number of vertices, the input to the encoder can be expressed in this reduced basis. For any configuration $\mathbf{u}$, we first project it using the continuous basis to obtain an $r$-dimensional feature $\mathbf{u}\,\mathbf{W}(\mathbf{x}_{0})^{\!\top}$, which is independent of mesh resolution. The encoder is then parameterized as a mapping $E:\mathbb{R}^{r}\!\rightarrow\!\mathbb{R}^{k},\;k<r$, and the latent coordinate is computed as $E(\mathbf{u}\,\mathbf{W}(\mathbf{x}_{0})^{\top})$.
We optimize the continuous model by minimizing the reconstruction loss over all training configurations and their sampled vertices:
(11) $\mathcal{L}_{\text{crom}}=\sum_{\mathbf{u}_{i}\in Q}\sum_{\mathbf{x}_{0}^{j}\in\mathcal{V}_{i}}\big\|\mathbf{W}(\mathbf{x}_{0}^{j})\,f\Big(E\big(\mathbf{u}_{i}\,\mathbf{W}(\mathbf{x}_{0}^{j})^{\!\top}\big)\Big)-\mathbf{u}_{i}(\mathbf{x}_{0}^{j})\big\|^{2},$
where $\mathcal{V}_{i}$ denotes the set of sampled rest positions (e.g., vertices) for configuration $\mathbf{u}_{i}$, and $\mathbf{u}_{i}(\mathbf{x}_{0}^{j})\in\mathbb{R}^{d}$ is the displacement at $\mathbf{x}_{0}^{j}$.
@@ 4.5.2. Continuous Model Reduction
Latent
dim. Reconstruction Error ($\downarrow$)
Vanilla AE Ours Linear Basis
Cute Dragon 8 $1.46\times 10^{-7}$ $3.62\times 10^{-7}$ $1.27\times 10^{-5}$
Xyz Dragon 8 $2.60\times 10^{-8}$ $7.90\times 10^{-8}$ $2.28\times 10^{-6}$
2D Bar 2 $7.51\times 10^{-7}$ $2.75\times 10^{-6}$ $9.36\times 10^{-6}$
Unicorn 8 $4.14\times 10^{-4}$ $2.22\times 10^{-4}$ $4.78\times 10^{-5}$
Hand 6 $5.11\times 10^{-8}$ $1.40\times 10^{-7}$ $3.00\times 10^{-7}$
Frog 8 $3.12\times 10^{-8}$ $8.08\times 10^{-8}$ $8.76\times 10^{-9}$
Armadillo 8 $1.46\times 10^{-7}$ $3.62\times 10^{-7}$ $1.27\times 10^{-5}$
Boat 3 $3.16\times 10^{-4}$ $1.70\times 10^{-3}$ $5.38\times 10^{-2}$
Whale 8 $3.49\times 10^{-8}$ $7.84\times 10^{-8}$ $8.33\times 10^{-7}$
Table 1. Reconstruction error comparison under identical latent dimensions. All methods in this table are evaluated using the same latent dimension to assess reconstruction capacity under a fixed-dimensional setting. In the full experimental pipeline, the latent dimension of the linear basis is increased as needed to match the reconstruction accuracy of nonlinear methods. Under this constraint, our method achieves reconstruction accuracy comparable to the vanilla autoencoder while outperforming the linear basis in most cases.
Unseen compression Reconstruction of Training Larger stretching Average Conv. time
Vanilla AE 22.38% 1.67% 4.80% 0.1729s
Linear Basis 13.51% 1.75% 3.31% 0.3258s
Ours 14.07% 1.69% 3.26% 0.1296s
Table 2. Average percentage difference from the ground-truth deformation and the convergence time, aggregated across three evaluation scenarios and over more than 50 test cases. Across these scenarios, our method consistently exhibits lower deviation than the vanilla autoencoder, while achieving error levels comparable to the linear basis, which similarly preserves the underlying deformation structure. In addition, our method converges fastest on average across the 50+ test cases.
@@ 5. Results
We evaluate our method across a range of settings. All networks are trained with Adam (Kingma and Ba, 2015) for 50k epochs at a learning rate of $10^{-4}$. For mesh-based models, the first and last layers are initialized using PCA. Unless otherwise specified, training data are generated with a full-space FEM solver using stable Neo-Hookean elasticity (Smith et al., 2018).
@@ 5.1. Generalization on Unseen Loading
We evaluate our model’s generalization capability under two out-of-distribution (OOD) scenarios, loading direction and loading magnitude, using three representative cases: opposite-direction loading on the Frog (Fig. 7), larger-magnitude loading on the Armadillo (Fig. 8), and increased-gravity loading on a 2D bar (Fig. 9). In all cases, models are trained within a limited loading regime and evaluated beyond the training distribution to assess robustness.
For the Frog (Fig. 7), the training data consist solely of compressive deformations generated by downward loading. At test time, we apply an equal-magnitude tensile load in the upward direction, implemented via fixed-point constraints. Starting from the rest state, we directly iterate under this unseen loading condition. We evaluate generalization by tracking the potential energy across Newton iterations, using energy decay as a measure of convergence.
@@
This setup explicitly probes the model’s ability to extrapolate beyond the training regime and recover physically consistent behavior under reversed loading conditions. The linear basis model reduction preserves a coherent, physically meaningful stretching pattern across iterations. The vanilla autoencoder (AE) remains numerically stable but fails to produce a plausible stretching deformation, often generating distorted shapes inconsistent with expected tensile behavior. In contrast, our symmetric odd-function formulation achieves shape quality comparable to the linear basis, faithfully capturing the intended stretch. Moreover, it converges to a low-energy configuration faster, highlighting its improved generalization under unseen loading directions.
@@
For the Armadillo (Fig. 8), the training dataset covers moderate loading magnitudes, while the test scenario applies substantially stronger forces that exceed the training range. This experiment evaluates the framework’s ability to generalize to high-strain regimes where nonlinear effects dominate. As in Fig. 7, we implement fixed-point constraints and iterate directly from the rest state via energy minimization. As shown in Fig. 8, our symmetric and convex formulation achieves markedly faster energy reduction than the linear baseline, while the vanilla AE plateaus at a higher energy level and fails to reach the minimal state. These results highlight our formulation’s efficiency and robustness when extrapolating to unseen, high-magnitude loading conditions.
This improved generalization also enables more stable simulation. For the 2D bar (Fig. 9), the training data are generated under normal gravity, under which all methods produce stable and physically plausible rollouts in the latent-space dynamics. When the gravity magnitude is increased by a factor of four—exceeding the training range—the same simulation procedure reveals markedly different behavior. Under this stronger loading, both the linear basis and our proposed method remain stable and preserve coherent deformation behavior, whereas the vanilla AE becomes unstable during the rollout, exhibiting non-physical deformations. This result indicates that, when extrapolating to larger-magnitude loading, enforcing structure in the latent mapping substantially improves robustness.
Figure 9. Latent-space dynamics under increased loading magnitude. Rollouts of a 2D bar under normal gravity (training regime) and under four-times increased gravity (out-of-distribution). While all methods exhibit stable dynamics under normal gravity, the vanilla autoencoder becomes unstable under stronger loading, producing non-physical behavior. In contrast, both the linear basis and the proposed structured formulation maintain stable and coherent deformation dynamics when extrapolating to larger-magnitude loading.
5.2. Performance Comparison on 50+ Evaluation Cases
Figure 10. Evaluation on a diverse set of over 50 test cases. The left panel shows the collection of shapes used for testing. The right panel illustrates the comparison setup: models are trained on a single deformation trajectory (top), and then evaluated under three settings: generalization to unseen compression, reconstruction of the training data, and generalization to larger stretching. This evaluation protocol is applied consistently across all test cases.
We also evaluate our method on a collection of 50+ test cases (Figure 10). The evaluation set includes multiple object instances, each subjected to several distinct deformation scenarios. For each object, the model is trained on a single deformation trajectory and then evaluated on that same loading as well as two additional, unseen loadings. These loadings are implemented using fixed-point constraints, and Newton’s method is used to minimize the elastic energy for a single static frame. Performance is measured by comparing the resulting deformed configurations against the corresponding ground-truth states.
Across all test cases, our method consistently achieves accurate matching on the two unseen deformations. It outperforms the vanilla autoencoder on both unseen cases and shows comparable performance to the linear basis on one case, while outperforming it on the other (Table 2). Although linear bases perform well for certain unseen deformations due to their fixed structured formulation, they require substantially higher reduced dimensions and consequently incur higher computational cost. In contrast, our method achieves comparable or better accuracy on unseen deformations using a more compact reduced representation, resulting in nearly a twofold reduction in computation time. For the training deformation, all methods accurately match the final configuration, with the vanilla autoencoder occasionally achieving lower error due to its unconstrained expressiveness on the observed data. Overall, our method provides a more favorable trade-off between accuracy, generalization robustness, and computational efficiency across deformation scenarios.
5.3. Comparison with Other Regularization Methods
Figure 11. Comparison of neural regularization methods on the hand model. All methods are trained on the same hand motion trajectory (top left) and evaluated on an unseen movement. Our method achieves the lowest final energy and reconstruction error while preserving natural and physically plausible deformations. In contrast, prior regularization approaches (Liu et al., 2022; Lyu et al., 2024) neither converge to the lowest energy state nor maintain deformation quality under extrapolation, resulting in higher error and visible artifacts.
We further compare our method with two existing regularization methods (Lyu et al., 2024; Liu et al., 2022). All methods are trained on the same deformation trajectory of the hand shape and evaluated by applying an unseen movement.
In this comparison (Fig. 11), our method achieves the lowest final energy and the smallest mean squared error (MSE) in the resulting deformation, outperforming both existing regularization baselines within the simulation setting. While both baseline methods aim to regularize nonlinear models via Lipschitz constraints, they differ in formulation and were originally developed for different tasks. In particular, the Lipschitz network (LipMLP) (Liu et al., 2022) was designed for neural SDF representation, where it performs well. However, when applied to our deformation task, LipMLP tends to produce less natural results and shows higher sensitivity to hyperparameter choices. This highlights the importance of tailoring regularization strategies to the structure and requirements of physical simulation.
The accelerated Lipschitz network (AccLip) (Lyu et al., 2024) did the regularization by incorporating Hessian-based energy information into the training loss, resulting in improved performance within the training distribution. However, it still shows noticeable deviations from the ground truth in certain regions—such as reduced bending in some fingers—and incurs substantially higher training costs, requiring approximately 22 hours under the same training setup due to explicit evaluation of Hessians with respect to the latent variables.
Under time-matched constraints (e.g., 20 minutes). While our method and AccLip both target physically informed latent-space models, they pursue different goals: AccLip focuses on accelerating convergence within the training regime, whereas our approach emphasizes generalization to unseen loadings. We view these directions as complementary and potentially synergistic.